### v1

# Characterizing Multi-Hunk Patches: Divergence, Proximity, and LLM Repair Challenges

Abstract
Multi-hunk bugs, where fixes span disjoint regions of code, are common in practice, yet remain underrepresented in automated repair. Existing techniques and benchmarks predominantly target single-hunk scenarios, overlooking the added complexity of coordinating semantically related changes across the codebase. In this work, we characterize Hunk4J, a dataset of multi-hunk patches derived from 372 real-world defects. We propose hunk divergence, a metric that quantifies the variation among edits in a patch by capturing lexical, structural, and file-level differences, while incorporating the number of hunks involved. We further define spatial proximity, a classification that models how hunks are spatially distributed across the program hierarchy. Our empirical study spanning six LLMs reveals that model success rates decline with increased divergence and spatial dispersion. Notably, when using the LLM alone, no model succeeds in the most dispersed Fragment class. These findings highlight a critical gap in LLM capabilities and motivate divergence-aware repair strategies.
I Introduction
Automated Program Repair (APR) has received significant attention from the research community [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. APR techniques have progressed from heuristic-based methods [19, 20, 21, 22] to deep learning approaches that learn repair patterns from large codebases [15, 16, 17, 10, 18], achieving promising results on single-hunk bugs [16, 17, 18, 10, 5, 4, 23, 24, 25, 26, 27]. Despite recent progress, most APR research remains focused on single-hunk scenarios [16, 24, 17, 18, 10], overlooking multi-hunk bugs, defects that require edits across disjoint code regions. These are both common in practice [28] and harder to fix due to their semantic interdependencies. While the problem has been acknowledged [28], only a few studies have directly addressed it [29, 28, 30, 31, 32, 33], and they have demonstrated limited effectiveness.
Recent advancements in large language models (LLMs) have shown promise in transforming several software engineering tasks, including program repair [34]. Trained on massive datasets comprising code and natural language, LLMs such as GPT-4 and LLaMA have demonstrated the ability to understand and generate code [35, 36, 37, 34]. While these models have already been employed to address single-hunk bugs with notable success, their potential to handle multi-hunk code repair remains underexplored.
Early studies on LLM-based multi-hunk repair include tools such as AlphaRepair [35], RepairAgent [38], and fine-tuning strategies [39, 40]. Although these approaches demonstrate that LLMs can extend beyond single-hunk scenarios, the challenges of reasoning about semantically interdependent edits distributed across the codebase remain unaddressed.
We identify four challenges limiting progress. First, dataset variation impedes consistency: prior studies evaluate on synthetic benchmarks such as LMDefects [41], algorithmically curated tasks (EvalRepair-C++/Java) [40], and real-world datasets such as Defects4J [42], often under varying assumptions, input formats, and test harnesses. Second, bug-type ambiguity obscures evaluation: multi-hunk bugs span a wide spectrum, from edits in non-contiguous regions within a single function to interdependent changes across multiple files. Yet prior work does not systematically characterize this heterogeneity, often conflating distinct forms of multi-location repair such as multi-line, multi-function, and multi-file edits. Third, reporting inconsistencies hinder meaningful attribution: existing techniques often report aggregate metrics that combine single- and multi-hunk results [40, 38], making it difficult to isolate and evaluate their effectiveness specifically on multi-hunk repairs. Fourth, existing evaluations lack patch complexity-aware metrics that quantify differences in hunks within a patch.
To address these challenges, we (1) propose hunk divergence, a metric that quantifies the internal variation within a multi-hunk patch by measuring lexical, structural, and file-level distance among hunks, while accounting for the number of hunks and the dispersion across files; (2) define spatial proximity as a classification scheme that assigns each patch to one of five categories—Nucleus, Cluster, Orbit, Sprawl, or Fragment—reflecting increasing levels of dispersion across the program hierarchy, including methods, files, and packages; (3) characterize multi-hunk patches using these abstractions to analyze hunk divergence and inter-hunk relationships; (4) develop Birch, a platform for standardized evaluation of LLM-based multi-hunk repair that supports reproducible comparisons across models and configurations; and (5) conduct a large-scale empirical study, examining the impact of retrieval strategies, context granularity, and feedback on repair success and cost under varying degrees of divergence and dispersion.
In this work, we make the following contributions:
  • We propose hunk divergence, a metric that quantifies intra-patch variation across lexical, structural, and file-level distance, and spatial proximity, a classification of multi-hunk patches based on their spatial layout in the codebase.
  • We introduce Hunk4J, a benchmark of 372 real-world multi-hunk bugs. We also present the first systematic characterization of multi-hunk bugs, focusing on their hunk diversity and proximity.
  • We present Birch, a platform for evaluating the capabilities of LLMs on multi-hunk bugs. Our empirical study spans 6 LLMs, various modes including 2 code representations, and 3 retrieval strategies, while also examining the role of contextual scope and incorporating feedback from compilation errors and test failures.
  • Our empirical findings show that model accuracy declines significantly as hunk divergence increases and spatial proximity decreases. All evaluated models without augmentation failed to repair any bugs in the most dispersed Fragment class, underscoring the challenges posed by highly scattered hunks.
II Characterization of Multi-hunk Patches
Although multi-hunk patches have been studied [28, 33, 38], their structural and spatial properties and how they affect automated repair remain poorly understood.
Definition 1 (Hunk).
A hunk is a block of consecutive edits applied to a specific region in the source code. We define each hunk $h_{i}$ as a tuple:
$h_{i}=(\text{loc}_{i},\text{content}_{i},\text{file}_{i},\text{method}_{i})$
where:
  • $\text{loc}_{i}$: line range of the edit,
  • $\text{content}_{i}$: token sequence derived from the raw diff (concatenated deletions and additions),
  • $\text{file}_{i}$: file path of the edit,
  • $\text{method}_{i}$: enclosing method identifier, or None if outside any method.
Definition 2 (Multi-Hunk Patch).
A multi-hunk patch $P$ is a set of $n\geq 2$ distinct hunks, denoted as $P=\{h_{1},h_{2},\dots,h_{n}\}$, where the hunks modify non-contiguous regions of source code. $P$ is an atomic change intended to address a single bug.
Our insight is that multi-hunk patches can vary significantly along several dimensions. In some cases, the hunks are nearly identical, such as repeated renamings or boilerplate edits, while in others, they involve semantically distinct changes that differ both lexically and structurally. These edits can also vary in proximity, from hunks located within the same method to those scattered across multiple files. This observation led us to ask: to what extent do hunks within a patch differ, and can this variation be meaningfully quantified?
We present the first in-depth characterization of multi-hunk patches by introducing two new metrics: (a) hunk divergence, which quantifies lexical, structural, and file-level dissimilarity among hunks, and (b) spatial proximity, which classifies the spatial layout of hunks across the code hierarchy.
II-A Hunk Divergence
We propose hunk divergence, a new metric that captures lexical, structural, and file-level differences among hunks in a patch.
Definition 3 (Pairwise Hunk Divergence).
We define the pairwise hunk divergence between two hunks $h_{i}$ and $h_{j}$ of $P$ through a combination of lexical, structural, and file-level separation distances:
$Div(h_{1},h_{2})=\frac{D_{\text{lex}}\cdot(D_{\text{ast}}+\gamma\cdot D_{\text{file}})}{1+\gamma}$
where:
  • $D_{\text{lex}}(h_{i},h_{j})$ quantifies the lexical distance between the code token sequences $\text{content}_{i}$ and $\text{content}_{j}$, where each token sequence is derived directly from the raw textual diff of the corresponding hunk.
  • $D_{\text{ast}}(h_{i},h_{j})$ measures the structural distance between the abstract syntax tree (AST) nodes of the two hunks.
  • $D_{\text{file}}(h_{i},h_{j})$ quantifies the file-level separation between the files $\texttt{file}_{i}$ and $\texttt{file}_{j}$ that contain the hunks $h_{i}$ and $h_{j}$, respectively.
  • $\gamma$ is a context-sensitive weighting factor that amplifies the contribution of inter-file separation to the overall divergence score, increasing the impact of file-based fragmentation when edits span distinct files in the repository.
By construction, if $D_{\text{lex}}$, $D_{\text{ast}}$, and $D_{\text{file}}$ are each normalized to the interval $[0,1]$, then the resulting pairwise divergence score satisfies $\text{Div}(h_{i},h_{j})\in[0,1]$.
To compute lexical distance ($D_{\text{lex}}$), we tokenize the diff exactly as it appears in the patch, preserving the ordering of added and deleted lines. For hunks involving both additions and deletions, we concatenate the deleted lines (before the change) and the added lines (after the change) prior to tokenization. In cases where a hunk contains only additions or only deletions, the token sequence consists solely of the respective lines as shown in the diff. We use BLEU [43] to quantify lexical similarity between token sequences of code fragments, following prior work [44, 45, 46, 47, 48, 17, 10, 49, 50].
$D_{\text{lex}}=1-\textsc{BLEU}(T_{1},T_{2})$
where $T_{1}$ and $T_{2}$ are the token sequences of hunks $h_{1}$ and $h_{2}$, respectively.
The goal of structural distance ($D_{\text{ast}}$) is to quantify how far apart two buggy hunks in the AST are when they appear in the same file. For this, we compute the shortest path between their corresponding nodes in the AST of the buggy (pre-patch) version of the file. Each hunk is associated with the AST node corresponding to the first statement that begins within its line range. For deletions, we extract this node from the pre-patch AST using the first deleted line in the hunk. Let $f=\text{file}(h_{1})=\text{file}(h_{2})$ be the shared file, and let $T_{f}$ denote the buggy file’s AST. Let $n_{1},n_{2}\in T_{f}$ be the AST nodes associated with $h_{1}$ and $h_{2}$, respectively. The structural distance is then defined as the normalized distance between these nodes, using the diameter of $T_{f}$ as a scaling factor.
To capture local structural variation between hunks, we normalize their AST distance using a natural logarithm-scaled formulation.
[…]
This formulation ensures numerical stability by avoiding division by zero when the AST is malformed or corresponds to an empty file. We apply log normalization to reduce the impact of large tree diameters, which can otherwise suppress meaningful structural differences between closely located nodes. If the buggy hunks occur in different files, we set $D_{\text{ast}}=1$. This reflects that structural comparison is undefined across separate ASTs, and maximum structural divergence is used in this case.
The goal of file divergence ($D_{\text{file}}$) is to quantify the degree of file-level separation between two hunks based on the layout of their enclosing file paths within the repository hierarchy. Let $\texttt{file}(h_{i})$ denote the fully qualified file path associated with hunk $h_{i}$, represented as an ordered sequence of directory segments. For brevity, we write $f_{i}=\texttt{file}(h_{i})$ and define $D_{\text{file}}:\mathcal{H}\times\mathcal{H}\rightarrow[0,1]$ as:
$D_{\text{file}}(h_{i},h_{j})=\begin{cases}0&\text{if }f_{i}=f_{j}\\ 1-\dfrac{|\texttt{LCP}(f_{i},f_{j})|}{\max\big{(}|f_{i}|,\ |f_{j}|\big{)}}&\text{otherwise}\end{cases}$
Here, $\texttt{LCP}(f_{i},f_{j})$ denotes the length of the longest common prefix between the directory sequences $f_{i}$ and $f_{j}$, the two distinct buggy files, and ||$\cdot$|| ⋅ | denotes the number of directory components in a given path. This formulation produces a normalized estimate of file-level distance: file paths that share a deeper hierarchical structure yield lower divergence values. This provides a continuous, repository hierarchy-aware quantification of inter-file separation that is both bounded and independent of path depth. It captures whether a pair of hunks is confined to the same file or spans across distinct files, serving as a proxy for the inter-file coordination effort required during multi-hunk repair.
Overall, our formulation of pairwise divergence $Div(h_{1},h_{2})$ ensures that the score remains bounded in $[0,1]$, while amplifying file-level separation when edits span multiple files. For same-file edits ($D_{\text{file}}=0$), the score is proportionally downweighted by $1+\gamma$, reflecting the reduced coordination effort relative to inter-file changes.
Input: Hunks $h_{1}$, $h_{2}$ with attributes: file, ast, code;
Output: Normalized divergence score $D(h_{1},h_{2})\in[0,1]$
1
$\gamma\leftarrow 1$
  // Scaling factor for file-level divergence
2
[…]
4
5if $h_{1}.\texttt{file}=h_{2}.\texttt{file}$ then
6       $f\leftarrow h_{1}.\texttt{file}$
7       $\mathcal{T}_{f}\leftarrow\textsc{ParseAST}(f)$
8       $treeDiameter\leftarrow\textsc{ComputeDiameter}(\mathcal{T}_{f})$
9
10      if $treeDiameter>0$ then
11             $n_{1}\leftarrow\textsc{LocateNode}(\mathcal{T}_{f},\ h_{1}.\texttt{loc})$
12
13            $n_{2}\leftarrow\textsc{LocateNode}(\mathcal{T}_{f},\ h_{2}.\texttt{loc})$
14
15            $astDist\leftarrow\textsc{ASTDistance}(n_{1},\ n_{2})$
16
17            $D_{\text{ast}}\leftarrow\ln(1+astDist)/\ln(1+treeDiameter)$
18
19       else
20             $D_{\text{ast}}\leftarrow 0$
21       end if
       $D_{\text{file}}\leftarrow 0$
        // No file-level separation: edits are in the same file
22
23 else
       $D_{\text{ast}}\leftarrow 1$
        // Maximum divergence across separate ASTs
24
25      $f_{1}\leftarrow h_{1}.\texttt{file}$, $f_{2}\leftarrow h_{2}.\texttt{file}$
26
27      $lcpLen\leftarrow\textsc{LCP}(f_{1},\ f_{2})$
28
29      $D_{\text{file}}\leftarrow 1-lcpLen/\max(|f_{1}|,\ |f_{2}|)$
30
      $\gamma\leftarrow 2$
        // Amplify inter-file separation weight
31
32 end if
33
[…]
35
[…]
37
38return $D(h_{1},h_{2})\leftarrow pairDiv/maxDiv$
Algorithm 1 Pairwise Divergence Score
[…]
Definition 4 (Hunk Divergence).
The hunk divergence of a multi-hunk patch $P$ quantifies the degree of internal variation among its constituent hunks. Let $P$ consist of $n$ hunks $\{h_{1},h_{2},\dots,h_{n}\}$, we define the hunk divergence of $P$ as:
$\text{Div}(P)=\ln(n)\cdot\left(\frac{2}{n(n-1)}\sum_{1\leq i<j\leq n}\text{Div}(h_{i},h_{j})\right)$
$\text{Div}(P)$ captures how lexically, structurally, and spatially dissimilar the hunks are pairwise and the coordination complexity introduced by the number of hunks in $P$.
We apply a natural logarithmic scaling factor $\ln(n)$ to reflect coordination complexity, increasing divergence for more fragmented patches while avoiding over-penalization when hunks are highly similar. As a result, the overall hunk divergence score satisfies $\text{Div}(P)\in[0,\ln(n)]$.
Consider the following examples: for a patch with $n=5$ hunks that are all identical, all pairwise distances are zero, yielding a hunk divergence score of $\text{Div}(P)=0$. At the other extreme, if all five hunks are maximally different (i.e., each pair has a divergence score of 1), then the average pairwise divergence is 1 and the resulting hunk divergence is $\text{Div}(P)=\ln(5)\approx 1.61$. In a more typical case, if the average pairwise divergence among the hunks is moderate (e.g., 0.5), the divergence score becomes $\text{Div}(P)=\ln(5)\cdot 0.5\approx 0.80$.
II-B Spatial Proximity
Spatial proximity classifies each multi-hunk patch into one of five categories—Nucleus, Cluster, Orbit, Sprawl, or Fragment—representing increasing levels of proximity dispersion across methods, files, and packages. Unlike hunk divergence, which assigns continuous quantitative scores based on hunk dissimilarity, this classification provides rule-based insights into the spatial distribution of hunks.
Definition 5 (Spatial Proximity Class).
Let $P=\{h_{1},\dots,h_{n}\}$ be a multi-hunk patch.
$\mathrm{SM}(P)$ $=\exists m\;\text{such that}\;\forall h_{i}\in P,\;m_{i}=m$
$\mathrm{SF}(P)$ $=\exists f\;\text{such that}\;\forall h_{i}\in P,\;f_{i}=f$
$\mathrm{SP}(P)$ $=\exists\mathcal{P}\;\text{such that}\;\forall h_{i}\in P,\;\mathcal{P}_{i}=\mathcal{P}$
$\mathrm{LCP}(P)$ $=\min_{i\neq j}\left|\text{LCP}(\mathcal{P}_{i},\mathcal{P}_{j})\right|$
These predicates assess different aspects of spatial proximity among the edits in a patch. Specifically:
  • $\mathrm{SM}(P)$: all hunks are in the same method.
  • $\mathrm{SF}(P)$: all hunks are in the same file.
  • $\mathrm{SP}(P)$: all hunks share the same package.
  • $\mathrm{LCP}(P)$: minimum shared directory depth across hunk pairs.
Then, the spatial proximity class C(P) is defined as:
$\mathcal{C}(P)=\begin{cases}\texttt{Nucleus}&\text{if }\mathrm{SF}(P)\land\mathrm{SM}(P)\\ \texttt{Cluster}&\text{if }\mathrm{SF}(P)\land\lnot\mathrm{SM}(P)\\ \texttt{Orbit}&\text{if }\lnot\mathrm{SF}(P)\land\mathrm{SP}(P)\\ \texttt{Sprawl}&\text{if }\lnot\mathrm{SF}(P)\land\lnot\mathrm{SP}(P)\land\mathrm{LCP}(P)>\lambda\\ \texttt{Fragment}&\text{if }\lnot\mathrm{SF}(P)\land\lnot\mathrm{SP}(P)\land\mathrm{LCP}(P)\leq\lambda\\ \end{cases}$
Here, $\lambda$ is a threshold on the directory depth used to distinguish structurally dispersed patches (Sprawl) from widely scattered ones (Fragment). This distinction is motivated by the intuition that when edits occur in directory paths with minimal shared hierarchy, the coordination burden for developers increases substantially.
II-C Hunk4J: A Dataset of Multi-Hunk Bugs
To enable principled evaluations, we introduce Hunk4J, a curated dataset of multi-hunk bugs derived from Defects4J [51]. Defects4J includes reproducible, test-suite backed bugs from open-source Java projects across diverse domains. We analyze all the 835 bugs in Defects4J [^1] We use Defects4J version 2.0.1 for our analysis. and identify those whose developer-written patches contain multi-hunk fixes (Definition 2). We found that 372 out of 835 bugs (44.6%) are multi-hunk. These 372 multi-hunk bugs form the basis of our Hunk4J for further analysis. Beyond extracting patch-level characteristics, we augment Hunk4J with natural language context for each bug. Specifically, we collect bug reports, titles, and summaries from official project repositories and issue trackers, including Apache JIRA, GitHub Issues, and SourceForge. This enriched metadata facilitates prompt construction and retrieval-augmented strategies for LLM-based program repair.
By combining patches validated by test suites with detailed hunk information (Definitions 1, 2), Hunk4J enables analysis of heterogeneity across lexical, structural, file-level divergence, and spatial distribution.
II-D Hunk Characterization
To analyze the complexity of multi-hunk bugs, we examine each patch in Hunk4J. Each patch is systematically analyzed based on the developer-written ground truth patch. We first identify the set of modified files by locating lines that begin with diff --git, which mark the inclusion of a file in the patch. Then, we extract distinct hunks within each file by detecting lines starting with @@, which indicate hunk boundaries. For single-hunk files, we further quantify the number of modified lines by counting those prefixed with either - (deletions) or + (additions). If a deletion is directly followed by an addition, we treat it as a single-line substitution rather than two independent changes. These steps facilitate a precise classification of hunks.
TABLE I: Multi-hunk patch file scope and hunk count.
File Scope Hunk Count Category Number of Bugs
Single-file Two hunks 140
Three hunks 55
Four or more hunks 49
Multi-file Two hunks 37
Three hunks 23
Four or more hunks 68
Total 372
We categorize multi-hunk bugs along two axes: the number of change regions (hunks) and the number of files touched in Table I. Single-file multi-hunk bugs dominate, with 244 cases where changes remain confined within a single file—this includes 140 two-hunks, 55 with three hunks, and 49 with four or more. These reflect intra-file repair scenarios, requiring local reasoning. In contrast, 128 bugs span multiple files, representing more fragmented, distributed repairs. Notably, 68 of these involve four or more hunks, indicating considerable dispersion across the codebase. For instance, multi-hunk patch $P$ for Jsoup_56 [52] touches 10 hunks across 5 files.
To complement our earlier categorization, Table II presents project-specific statistics capturing how multi-hunk bugs manifest across different codebases. In 15 out of 17 projects, the median number of hunks per bug is at most 3, suggesting that the majority of bugs involve relatively localized changes. However, outliers exist, Jsoup_87 [53] has 47 hunks. In contrast, Chart_4 [54] involves a multi-hunk fix with only two hunks confined to a single method, which is simpler to reason about. Across all 17 projects, the median number of files modified per bug does not exceed 2; nonetheless, notable exceptions are present. For instance, JacksonDatabind_103 [55] spans 26 hunks across 16 files, modifying exception handling across the codebase. Conversely, certain projects such as Lang, Collections, and JacksonXml contain no multi-file bugs.
TABLE II: Hunk4J dataset: hunk and file statistics
Project Bugs Hunks per Bug Files per Bug
Min Median Mean Max Min Median Mean Max
Chart 10 2 2.50 3.00 6 1 1.00 1.20 2
Cli 17 2 2.00 3.47 10 1 1.00 1.71 4
Closure 82 2 3.00 4.06 21 1 1.00 1.63 6
Codec 7 2 3.00 3.57 7 1 2.00 2.43 6
Collections 1 3 3.00 3.00 3 1 1.00 1.00 1
Compress 18 2 3.00 3.89 14 1 1.00 1.61 7
Csv 4 2 2.00 2.50 4 1 1.00 1.25 2
Gson 7 2 2.00 4.00 12 1 1.00 1.43 3
JacksonCore 12 2 3.50 4.33 8 1 2.00 1.75 3
JacksonDatabind 55 2 3.00 4.58 26 1 1.00 1.87 16
JacksonXml 3 2 2.00 2.67 4 1 1.00 1.00 1
Jsoup 37 2 3.00 4.73 47 1 1.00 1.89 5
JxPath 14 2 4.00 4.50 11 1 2.00 1.93 5
Lang 25 2 2.00 2.72 7 1 1.00 1.00 1
Math 50 2 2.00 2.80 12 1 1.00 1.26 7
Mockito 20 2 3.00 4.00 20 1 1.00 1.40 5
Time 10 2 3.00 4.00 10 1 1.00 1.50 3
Total 372 2 3.00 3.86 47 1 1.00 1.59 16
Figure 1 presents the distribution of divergence for Hunk4J. Lexical distance ($D_{\text{lex}}$) (Md = 0.94, Mean = 0.82) is sharply skewed toward the upper bound, indicating that most hunk pairs are lexically dissimilar. Structural distance ($D_{\text{ast}}$) (Md = 0.6, Mean = 0.62) shows broader dispersion, reflecting moderate variation in the AST of edits. File-level distance for multi-file patches ($D_{\text{file}}$) exhibits a moderate central tendency (Md = 0.250, Mean = 0.251), indicating that while many multi-hunk patches involve hunks within closely related files, a non-trivial portion span more distinct locations in the file hierarchy.
Overall hunk divergence in Hunk4J ranges from a minimum of 0.00 to a maximum of 1.60, with a median of 0.42, a mean of 0.47. Notably, 25% of multi-hunk patches exhibit divergence values greater than 0.624. The distribution reflects heterogeneity across lexical content, code structure, and file-level separation. Importantly, the divergence landscape is far from trivial, refuting any assumption of homogeneity.
(a) Lexical
(b) Structural
(c) File
(d) Overall
Figure 1: Distribution of different divergence components and the overall hunk divergence.
We also characterize the multi-hunk patches according to our spatial proximity classification (Definition 5). We instantiate the proximity threshold $\lambda$ based on the empirical distribution of package path depths in Hunk4J. The median package depth across all buggy files is 6; we set $\lambda=3$, corresponding to half this value, to conservatively separate loosely grouped edits from those that are structurally dispersed across the codebase.
Table III categorizes multi-hunk patches based on their spatial proximity, revealing a distinct skew toward more localized fixes. The Cluster class, comprising 185 bugs, is the most prevalent and includes patches where multiple hunks occur within the same file but in different methods. This is followed by the Orbit class (67 bugs), which captures cases where hunks span multiple files within the same package, and the Nucleus class (59 bugs), where all hunks are confined to a single method. Mockito_6 [56] from the Cluster class contains 20 hunks, all confined to a single file yet distributed across multiple methods, demonstrating file-level cohesion despite method-level dispersion. In contrast, the Sprawl (50 bugs) and Fragment (11 bugs) classes represent increasingly dispersed patches. The Fragment category consists of patches where hunks are scattered across unrelated packages. An example from Fragment class is Closure_144 [57], in which the hunks are scattered across unrelated packages, highlighting the low spatial locality. The average hunk divergence increases monotonically with spatial dispersion. Fragment patches, though rare, exhibit the highest divergence (0.7372), underscoring their complexity.
TABLE III: Spatial proximity class and hunk divergence.
Proximity Class Number of Bugs Mean Hunk Divergence
Nucleus 59 0.2548
Cluster 185 0.4280
Orbit 67 0.5628
Sprawl 50 0.6718
Fragment 11 0.7372
III Empirical Evaluation
Our study addresses the following research questions:
  • RQ1: How well do LLMs perform on multi-hunk repair tasks?
  • RQ2: To what extent do augmented prompting techniques (RAG, context, feedback) improve repair accuracy?
  • RQ3: How do hunk characteristics such as divergence and spatial proximity influence the repair success of LLM-based techniques?
  • RQ4: Are certain types of divergence and spatial proximity more predictive of challenges in LLM repair?
III-A Benchmarking Framework: Birch
To enable reproducible evaluations across different LLMs, we present Birch, a benchmarking framework for multi-hunk program repair. It offers an end-to-end pipeline, from input preparation to patch validation. The process begins by checking out the buggy version of the program from Hunk4J and setting up a workspace. For each bug, Birch extracts the buggy code, relevant test cases, and associated metadata. It then constructs an input prompt that includes the buggy code, the precise bug location, a natural language description of the issue, and the failing test cases with their corresponding error messages. Once the prompt is prepared, it is submitted to a selected LLM. The patch is parsed and applied to the buggy code. The code is then compiled and tested using the test suite. Compilation errors and failing test cases are logged.
Modes. Birch supports assessing vanilla LLMs but also provides various optional modes for enhancing LLMs. The retrieval-based example selection mode supports both sparse and dense retrieval techniques. The enclosing context selection mode extracts context at varying granularities, including the method, class, and file levels. The feedback loop mode facilitates iterative patch refinement: if a generated patch fails to compile, the corresponding compilation error message is appended to the subsequent prompt. If the patch compiles but fails to pass the test suite, detailed test failure information is incorporated to guide further refinement attempts. If the code compiles and passes all tests, the patch is marked as successful. All evaluation artifacts, including prompts, generated code, diffs, and test logs, are persisted for inspection and analysis.
Implementation. Birch is implemented in Python and supports both proprietary models and open-source models via Ollama [58]. We use litellm [59] to interface with LLM APIs across providers. Prompts are templated using TOML, enabling integration of new prompt variants. We employ JavaParser to parse ASTs, which are used to compute node-level distances between hunks, derive the tree diameter, and extract enclosing scopes such as methods and classes. For sparse retrieval, we use BM25; for dense retrieval, we employ text-embedding-text-small-3 or Mini-LM, storing embeddings in a FAISS [60] index for nearest‐neighbor lookup.
III-B Models
This study conducts a comparative evaluation of six models, selected to represent both proprietary and open-source variants with diverse architectural characteristics. The proprietary models include Gemini-2.5-Flash (1M token context window) from Google, evaluated in non-reasoning mode via the official Gemini API, and Nova-Pro (300K tokens) from Amazon. From OpenAI, we evaluate both a reasoning-capable model, o4-mini (200K tokens), and a non-reasoning variant, GPT-4.1 (1M tokens), accessed directly via the OpenAI API. Parameter counts for proprietary models are not publicly disclosed. The open-source models comprise mistral-large-2407 (128K tokens, 123B parameters) from Mistral AI and LLaMA3.3 (128K tokens, 70B parameters) from Meta. All models, except those from OpenAI and Google, were evaluated using full-parameter inference endpoints through Amazon Bedrock [61]. Our goal is not to perform exhaustive model benchmarking. Rather, we aim to assess the generalizability of our findings across LLMs with varying architectures.
III-C Evaluation Metrics
We assess patch correctness based on whether the generated code parses, compiles, and passes all tests provided by the benchmark suite. Following prior work [62], we report Plausible@1, defined as the fraction of bugs for which the top-1 generated plausible patch passes the complete test suite:
[…]
where $\mathcal{B}$ is the set of bugs and $\texttt{patch}_{1}^{b}$ is the first generated patch for bug $b$. A patch is counted as plausible if it compiles and passes all test cases, including the originally failing one.
We generate one patch per bug using standard decoding settings (e.g., temperature = 0.0) for all models except o4-mini, whose API has a fixed temperature of 1.0. While decoding is configured to be deterministic at the API level, generation remains non-deterministic in practice due to low-level system factors such as thread scheduling, memory layout, and backend randomness [63].
TABLE IV: Multi-hunk bug repair with LLMs
Type Model Pass Test Fail Comp Fail Plausible@1 Cost
Open Source Llama3.3 46 73 253 12.37% \$1.43
mistral-2407 45 106 221 12.10% \$4.08
Propri- etary nova-pro 33 82 257 8.87% \$2.30
Gemini 2.5 Flash 50 32 290 13.44% \$0.37
GPT-4.1 82 78 212 22.04% \$6.37
o4-mini 100 57 215 26.88% \$3.30
III-D Effectiveness of LLMs on Multi-Hunk Repair (RQ1)
To assess LLM effectiveness on multi-hunk repair, we ran all 372 bugs through the six models. Table IV summarizes the results, including the number of fully repaired bugs (Pass), test failures, compilation failures, plausible patch rate (Plausible@1), and estimated cost per run. Overall, the best-performing model was o4-mini, which successfully repaired 100 of the 372 bugs (26.88%), followed by gpt-4.1 with 82 successful repairs (22.04%). Among open-source models, LLaMA3.3 achieved the highest accuracy, repairing 46 bugs (12.37%), followed closely by mistral-2407, which repaired 45 bugs (12.10%).
Across the six LLMs, a total of $127$ unique bugs are resolved. Among the two OpenAI models evaluated, we selected O4-Mini for further analysis due to its higher accuracy compared to GPT-4.1. We then assess individual effectiveness and collective coverage using the following five models: o4-mini, Gemini-2.5-Flash, Nova-Pro, Mistral-2407, and LLaMA3.3. Specifically, we conduct an intersection analysis on the sets of bugs each model successfully resolves. Together, these five models generate plausible patches for $116$ unique bugs. This finding underscores that substantial coverage in multi-hunk bug repair can be achieved through a carefully selected ensemble of models. Figure 2 presents a Venn diagram showing the overlap of Plausible@1 patches among the selected models. Although all five models solve a shared pool of $16$ bugs, each contributes additional unique fixes, most notably o4-mini, which accounts for $39$ unique repaired bugs.
When considering cost, model accuracy varies significantly relative to inference expense. Gemini-2.5-Flash achieves a Plausible@1 of $13.44\%$ at a total cost of just \$0.37, the lowest among all evaluated models. In contrast, Mistral-Large-2407, an open-source model, incurs a substantially higher cost of \$4.08 while yielding a comparable Plausible@1 of $12.10\%$. This disparity highlights that increased inference cost does not necessarily translate to higher repair accuracy.
Overall, despite recent advances, LLMs still struggle with multi-hunk bug repair, as even the best model fails to fix over 70% of cases. The higher accuracy of o4-mini highlights the benefit of using reasoning-capable models with greater test-time compute compared to non-reasoning models. However, the limited repair accuracy even in the best-performing model suggests that increasing test-time compute alone may not be sufficient, particularly for complex repair tasks [64].
Figure 2: Overlap of fixes across model families
III-E Impact of Augmented Prompting (RQ2)
To assess the impact of different prompt augmentation strategies on LLM-based program repair, we evaluate three aspects: (a) retrieval-augmented example selection, (b) contextual granularity, and (c) feedback-based refinement. Given the high inference cost of large-scale evaluation across all prompt augmentation strategies, we limit this analysis to two representative models: o4-mini, the most accurate proprietary model, and LLaMA3.3, the most effective open-source model in our evaluation.
We examine the effectiveness of retrieval-based prompting by comparing different ways of representing and retrieving similar repair examples. Each buggy example is encoded either as code tokens or as AST elements, capturing lexical and structural perspectives, respectively. For each representation, we apply one sparse retrieval method and two dense embedding-based methods to identify semantically similar past patches as examples. By pairing the two code representations with these three retrieval techniques, we construct six example-selection strategies, each retrieving the top-5 most similar patches to include in the prompt. The example patches used in retrieval-augmented prompting are selected from multi-hunk patches in the Hunk4J dataset that were fixed prior to the target bug—excluding the target bug itself—based on similarity in token or AST representations, depending on the retrieval configuration.
Table V reports the effectiveness of various prompt augmentation strategies in terms of accuracy, failure outcomes, and inference cost on o4-mini and LLaMA3.3. The results show that token-level representations paired with dense retrieval outperform both AST-based code views and the sparse retrieval (BM-25) technique. For o4-mini, the best configuration—token+text-embed—achieves a Plausible@1 of 31.18%, surpassing the AST+text-embed variant (23.66%) by 7.52 percentage points. Among retrieval strategies, text-embed consistently yields higher accuracy than MiniLM and BM25 across both token and AST inputs. Similar trends are observed for LLaMA3.3, where token+text-embed reaches 20.16% Plausible@1, compared to 11.02% with AST+text-embed and 17.20% with token+BM25. This improvement can be attributed to how dense retrieval operates. Unlike sparse retrieval (BM25), which relies on token overlap and term frequency, dense retrieval (text-embed) encodes both the query and candidates into dense vectors using a neural embedding model. The embedding model was trained on token sequences, not structured ASTs. Consequently, similarity search is more accurate when both the query and candidates are represented as token sequences.
TABLE V: Multi-hunk bug repair with prompting strategies
Strategy LLM Prompt Mode Pass Test Fail Comp Fail Plausible @1 Cost
RAG (Example Selection) Llama3.3 AST + BM25 64 76 232 17.20% \$3.48
AST + MiniLM 42 96 234 11.29% \$9.09
AST + text-embed 41 94 237 11.02% \$7.94
Token + BM25 70 75 227 18.82% \$4.77
Token + MiniLM 69 75 228 18.55% \$3.75
Token + text-embed 75 77 220 20.16% \$4.00
o4-mini AST + BM25 106 66 200 28.49% \$6.53
AST + MiniLM 91 74 207 24.46% \$14.90
AST + text-embed 88 79 205 23.66% \$13.25
Token + BM25 111 57 204 29.84% \$8.46
Token + MiniLM 114 57 201 30.65% \$8.38
Token + text-embed 116 59 197 31.18% \$7.26
Context Granularity Llama3.3 Method 46 73 253 12.37% \$1.43
Class 43 113 216 11.56% \$3.80
File 58 132 182 15.59% \$4.03
o4-mini Method 100 57 215 26.88% \$3.30
Class 88 146 138 23.66% \$35.47
File 105 186 81 28.30% \$43.35
Feedback Loop Llama3.3 Method + Code + text-embed 91 80 201 24.46% \$7.64
o4-mini Method + Code + text-embed 133 34 205 35.75% \$14.63
TABLE VI: Hunk divergence for fixed and unfixed bugs, and spatial proximity distribution (fixed) across model families
Model Hunk Divergence (Fixed) Hunk Divergence (Unfixed) Spatial Proximity (Fixed)
Median Mean Max Std Dev Median Mean Max Std Dev Nucleus Cluster Orbit Sprawl Fragment
mistral-2407 0.27 0.27 0.65 0.19 0.46 0.49 1.60 0.30 20.34 11.35 5.97 16.0 0.0
Llama3.3 0.26 0.27 0.65 0.17 0.47 0.50 1.60 0.30 27.12 9.73 11.94 8.0 0.0
nova-pro 0.24 0.28 0.65 0.19 0.45 0.49 1.60 0.30 11.86 8.11 7.46 12.0 0.0
Gemini 2.5 0.26 0.26 0.65 0.17 0.47 0.50 1.60 0.30 18.64 12.43 14.93 12.0 0.0
o4-mini 0.29 0.28 0.70 0.18 0.50 0.53 1.60 0.30 49.15 23.78 23.88 22.0 0.0
In addition to example selection, we study the effect of contextual granularity by varying the scope of code included in the prompt. We compare three levels of enclosing context: method, class, and file. The results for enclosing context show that wider context generally improves repair success. For o4-mini, Plausible@1 increases from 26.88% (method) to 28.30% (file), with a drop at the class level (23.66%), as class-level prompting often result in incomplete code generation. The reduced effectiveness at the class level can be attributed to LLMs interpreting the class snippet as an incomplete code fragment, often defaulting to patching only individual methods or lower-level constructs. In contrast, file-level prompts, which include import statements and top-level comments, provide clearer semantic cues, prompting the model to generate complete code. LLaMA3.3 shows a similar pattern: file-level prompting yields 15.59%, compared to 12.37% (method) and 11.56% (class). However, these gains come at a significant cost. For o4-mini, file-level prompts cost \$43.35, 13 times more than method-level (\$3.30), which achieves nearly comparable accuracy.
We further examine feedback-based refinement, which augments prompting with dynamic execution information derived from compilation failures or test failures, following prior work [65, 66, 67]. To limit cost, the feedback prompt uses the method-level context and the best retrieval setup (token+text-embed). Despite this conservative design, feedback yields the highest accuracy across all settings. o4-mini achieves 35.75% Plausible@1 and LLaMA3.3 achieves 24.46%, both surpassing their respective static prompt baselines. Notably, the accuracy of o4-mini with feedback surpasses that of the static prompt using file-level context ($28.30\%$), while costing less than half as much (\$14.63 vs. \$43.35). These results indicate that feedback, driven by runtime information from failed tests or execution traces, can guide the model more effectively than simply expanding the prompt context. Moreover, feedback mitigates the limitations of narrow context by incorporating dynamic execution traces, which provide concrete, failure-specific guidance that static prompts lack.
Although prompting augmentations such as feedback and retrieval improve repair accuracy over vanilla LLMs, even with the best configurations, 64% of multi-hunk bugs still remain unfixed.
III-F Repair Outcomes by Hunk Divergence and Proximity (RQ3)
TABLE VII: Hunk divergence for fixed and unfixed bugs, and spatial proximity (fixed) across best prompt strategies
Model Technique Variant Divergence Score (Fixed) Divergence Score (Unfixed) Spatial Proximity (Fixed)
Median Mean Max Std Dev Median Mean Max Std Dev Nucleus Cluster Orbit Sprawl Fragment
o4-mini Token + text-embed 0.34 0.31 0.96 0.21 0.50 0.54 1.60 0.30 54.24 30.81 22.39 20.00 18.18
File Context 0.39 0.40 1.08 0.23 0.46 0.49 1.60 0.31 37.29 33.51 13.43 22.00 9.09
Feedback Loop 0.31 0.30 0.96 0.20 0.53 0.56 1.60 0.30 64.41 33.51 28.36 22.00 27.27
Llama3.3 Token + text-embed 0.28 0.28 0.78 0.21 0.47 0.52 1.60 0.29 33.90 18.92 17.91 14.00 9.09
File Context 0.36 0.36 1.08 0.23 0.45 0.49 1.60 0.30 18.64 17.30 7.46 20.00 0.00
Feedback Loop 0.30 0.29 0.96 0.20 0.49 0.52 1.60 0.30 42.37 21.62 25.37 14.00 18.18
To examine the conditions under which LLM-based program repair is most effective, we measure descriptive statistics of hunk divergence and spatial class-wise accuracy percentages. These metrics are derived from patches produced by the five models selected in RQ1, each representing a different model family.
Table VI reports the central tendencies of hunk divergence for both successfully and unsuccessfully fixed bugs. Additionally, it presents the proportion of bugs correctly fixed within each spatial proximity class.
Successful repairs are consistently associated with low divergence across all five models (mean 0.26absent0.26$\approx$ 0.26≈ 0.260.280.280.280.28; median 0.24absent0.24$\approx$ 0.24≈ 0.240.290.290.290.29). Variance remains narrow (standard deviation $<0.20$), indicating high semantic alignment across generated patches. In contrast, unfixed bugs exhibit substantially higher hunk divergence (mean: 0.490.490.490.490.530.530.530.53; median: 0.450.450.450.450.530.530.530.53), along with broader variance and higher maxima (up to $1.60$). These results reflect the increased difficulty LLMs face in coordinating complex, semantically diverse edits across scattered code regions.
Table VII reports our analysis across the best-performing prompting configurations, i.e., (1) Token + text-embed, (2) File Context, and (3) Feedback Loop, applied to o4-mini and Llama3.3. Across all strategies, fixed bugs consistently exhibit lower divergence than unfixed bugs. The mean divergence for fixed bugs ranges from $0.28$ to $0.40$ (median: 0.280.280.280.280.390.390.390.39), while unfixed bugs display substantially higher scores (mean: 0.490.490.490.490.560.560.560.56; median: 0.450.450.450.450.530.530.530.53), indicating that greater heterogeneity within hunks correlates with repair failure. Among the evaluated strategies, Feedback Loop achieves the lowest divergence values for fixed bugs in both models (mean: 0.290.290.290.290.300.300.300.30; median: 0.300.300.300.300.310.310.310.31), and the highest percentage of repairs in the Nucleus and Cluster classes. This pattern suggests improved effectiveness on bugs involving spatially localized edits. These results demonstrate that prompting strategy significantly influences both semantic coherence and spatial locality of successful patches. Notably, the Feedback Loop configuration demonstrates the highest effectiveness in achieving low-divergence repairs within high-proximity contexts, thereby facilitating more precise and coordinated multi-hunk program repair.
The association between spatial proximity and repair success is first evident in Table VI. Repair success is highest in the Nucleus and Cluster classes, where hunks are in closer proximity, while substantially lower fix rates are observed in the Sprawl and Fragment categories, with the latter exhibiting no successful repairs. This trend persists across prompt variants, as shown in Table VII. Among the evaluated strategies, Feedback Loop consistently achieves superior accuracy in close proximity. For example, o4-mini with Feedback Loop repairs 64.41% of bugs in the Nucleus class and 33.51% in Cluster, compared to only 22.00% and 27.27% in Sprawl and Fragment, respectively. A comparable pattern is observed for Llama3.3.
Figure 3: Upset-style plot showing bugs correctly repaired by the selected models, grouped by spatial proximity class. Bars indicate intersections among models; color segments denote proximity categories.
To further investigate whether the spatial proximity of hunks influences the success of LLM-based repair, we analyze the subset of bugs that were correctly fixed by at least one top-performing model from five distinct LLM families. Out of 372 multi-hunk bugs in Hunk4J, a total of 116 were successfully repaired by at least one model. Figure 3 presents an UpSet-style visualization of these 116 fixed bugs, grouped by spatial proximity class and by the specific combinations of models that produced successful patches. Each bar along the horizontal axis represents a unique combination of models (denoted by filled circles in the matrix below), and bars are ordered from left (representing the fewest bugs) to right (representing the most bugs). Within each bar, segments are color-coded by proximity class, Nucleus (dark blue), Cluster (light blue), Orbit (orange), Sprawl (peach), and Fragment (green) with each segment labeled by the exact bug count it represents. The total number of bugs in each model intersection is annotated above the corresponding bar. The lower panel presents a matrix of filled circles and vertical connectors: each row denotes one of the 5 models (from top to bottom: Mistral-2407, LLaMA3.3, gemini-2.5-flash, Nova-Pro, o4-mini), and each column corresponds to the same model combination that generated the bar above. A filled circle indicates that the model participated in that combination; connectors link the participating models.
Two trends are evident. First, the majority of successful repairs fall into the Nucleus and Cluster categories, which represent cases where the relevant code elements are tightly co-located. These proximity classes dominate across both shared and model-specific intersections. Second, cases in the Orbit and Sprawl categories, where relevant context is more dispersed, are less frequently solved. The LLM alone was unable to fix any bug in the Fragment category, which reflects extreme dispersion or disconnection of relevant context (Table VI). The largest set of bugs (39 bugs) is uniquely attributable to o4-mini, while the second-largest (16 bugs) was solved by all five models. Smaller intersections reveal model-specific capabilities or divergences, particularly in handling less proximal context.
Collectively, these findings suggest that successful LLM-based repairs for multi-hunk bugs are characterized by (1) semantically coherent edits with low hunk divergence and (2) spatially localized dependencies with high proximity.
III-G Divergence, Proximity, and Repair Difficulty (RQ4)
We now investigate whether hunk divergence and spatial proximity function as predictive indicators of LLM repair difficulty, i.e., the likelihood that a model fails to synthesize a correct patch. This analysis considers two dimensions of multi-hunk patches: the degree of intra-patch variability (hunk divergence) and the extent of spatial dispersion across the program hierarchy (spatial proximity).
To assess this, we first partitioned the data for the Top-5 LLMs into two independent samples, fixed and unfixed bugs, and applied a two-sample Wilcoxon rank-sum test to their divergence scores. Across all five top-performing models, the divergence distributions for fixed and unfixed bugs differ significantly. All Wilcoxon tests yield $p$-values below $10^{-4}$, with three models reporting $p<10^{-7}$ and the largest $p$-value at $2.6\times 10^{-5}$. These results confirm that higher divergence is strongly associated with failure to generate a correct patch. We also computed Cliff’s Delta to assess the effect sizes. Under LLM-only and feedback-loop prompting, the delta values for o4-mini were $0.533$ and $0.554$, respectively. Both values exceed the threshold of $0.474$, indicating large effect sizes, that is, substantial differences in hunk divergence between successfully and unsuccessfully repaired bugs. The consistent statistical significance observed across models suggests that hunk divergence may serve as a meaningful indicator of patch complexity and the effectiveness of LLM-based repair.
We observe similar trends when comparing augmented prompting configurations (Table VII). Among the six prompting variants, four yield $p$-values below $10^{-7}$, one yields $p=3.8\times 10^{-3}$, and all remain below the conventional threshold of $0.05$. This demonstrates that the distinction between fixed and unfixed bugs based on divergence holds across different prompt strategies, further validating its predictive value.
Figure 4 presents faceted violin plots showing the distribution of average hunk divergence scores for each model, separated by outcome. Each violin illustrates the density of divergence values, while the embedded boxplot indicates the median and interquartile range, capturing the central 50% of the data. Fixed bugs are consistently associated with lower divergence scores across all models. This clear and consistent pattern demonstrates that hunk divergence is a discriminative signal for reasoning about multi-hunk bug complexity and the success of LLM-based repair.
Figure 4: Faceted violin plots of average hunk divergence, grouped by model and split by outcome.
In addition to divergence, spatial proximity also emerges as a strong categorical predictor. As shown in Figure 3, repair success declines consistently from Nucleus to Sprawl, with the Fragment class remaining entirely unsolved across all models. This trend suggests that as fault-relevant context becomes more dispersed, LLMs are increasingly unable to synthesize correct patches. The complete absence of fixes in the Fragment category indicates a hard boundary for current model capabilities in handling non-local, scattered dependencies.
Our analysis shows that LLMs exhibit a significant drop in repair accuracy as hunk divergence increases and spatial proximity decreases. Bugs with high intra-patch variation or dispersed hunks across methods, files, or packages remain largely unrepaired. These findings underscore the need for models and prompts that explicitly account for lexical and structural dispersion in multi-hunk patches.
IV Threats to Validity
Hunk4J is derived exclusively from Java projects, which may limit generalizability to other languages. Although these bugs are real-world and diverse, their distribution reflects the design and testing philosophy of Java projects, which may bias the scope of repair challenges.
We use the Plausible@1 metric to reduce inference cost, which may understate model accuracy in iterative settings. Decoding parameters and prompt templates are fixed across models to ensure fairness, potentially favoring some architectures. However, our aim is comparative analysis under uniform conditions and not model-specific optimization.
Our evaluation assumes perfect fault localization, following prior LLM-based repair studies to isolate patch generation quality [35, 41, 33, 38, 34, 17, 10, 18]. While this sets an upper bound on accuracy, real-world scenarios often involve imperfect or incomplete localization. Notably, even with perfect localization, LLMs struggle with multi-hunk repairs, failing on all Fragment-class bugs (Table VI), highlighting challenges in semantic coordination beyond localization.
LLMs are accessed via third-party APIs, and execution can vary due to backend updates or system-level nondeterminism. Although decoding is configured to be deterministic (temperature = 0), minor variations may still occur. We mitigate this by logging all inputs, outputs, and execution metadata for reproducibility.
Our implementation uses JavaParser to extract ASTs and compute divergence metrics such as node-level distance and tree diameter. Alternative parsers (such as javalang) may yield slight variations due to differences in AST tree construction, though all models are evaluated under the consistent criteria.
To support reproducibility, we release Birch, all scripts, prompts, and the Hunk4J dataset [68].
V Related Work
Most APR benchmarks offer limited support for multi-hunk bugs. While Defects4J [51], Bugs.jar [69], BugsInPy [70], and LMDefects [41] include them, others offer limited coverage [71, 72, 73] or focus on single-hunk edits [74, 75, 76]. Existing multi-hunk LLM repair studies rely on synthetic [41], function-level [42], or algorithmic [40] benchmarks, limiting generalizability. We introduce Hunk4J, a real-world, multi-hunk benchmark derived from Defects4J, augmented with divergence metrics for evaluating LLM-based repair.
Prior studies have applied change metrics for fault prediction. Graves et al. [77] used coarse-grained measures such as file and modification counts; D’Ambros et al. [78] highlighted the predictive power of change coupling and scatter; and Ferzund et al. [79] proposed hunk-level metrics. In contrast, we quantify semantic and structural heterogeneity in multi-hunk patches to assess LLM-based repair effectiveness.
Prior work on multi-hunk repair with LLMs [41, 42, 40, 38, 39, 80, 35] exhibits wide variation in datasets, prompting strategies, and evaluation protocols, often reporting aggregate results without isolating multi-hunk accuracy. While prior studies have examined patch properties such as hunk divisibility [80], they do not account for spatial dispersion or intra-patch heterogeneity. We address this limitation by introducing Hunk Divergence and Spatial Proximity, two metrics that capture lexical, structural, and hierarchical relationships between hunks to enable complexity-aware analysis.
VI Conclusion
We introduce the first characterization of multi-hunk bugs through the lens of intra-patch divergence. Our metric, hunk divergence, captures lexical, structural, and file-separation dissimilarity between hunks, while spatial proximity classifies their distribution across the codebase. Our findings reveal that multi-hunk bugs exhibit substantial variability in divergence, with no single model achieving consistently high accuracy across this spectrum. Collectively, these contributions lay the foundation for developing divergence-aware techniques capable of addressing the unique challenges posed by multi-hunk bugs. In future work, we aim to develop a divergence-aware model selector that routes bugs to the most capable and cost-efficient LLM based on predicted repair complexity.